What is the total force acting on the outside of the spherical submarine of diameter 2.0 m when it is 1000 m below the surface of the ocean? The pressure inside is 1.0 ATM and the density of seawater is 1020 kg / m3
Answer #1

Pressure inside the submarine = 1 atm

pressure outside the submarine = atmospheric preesure ( 1 atm) + pressure due to water column at that depth (1000m)

therefore, net pressure on the submarine = pressure due to water column

= h * d * g ; where h is the depth of water , d is the density of sea water , g is the acceleration due to gravity

= 1000 * 1020 * 9.8

= 9996000 Pascals

using , Pressure = Force / Area

9996000 = Force / (4*pi*r^2)

Force = 9996000 * 4 * pi * (1)^2

Force = 125613440.7 Newton

Please ask your doubts or queries in the comment section below.

Please kindly upvote if you are satisfied with the solution.

Thank you.